### v2

@@ Abstract
Quasi-free scattering of atomic nuclei away from the stability line has reached several milestones over the past decade. The advent of gamma, charged-particles, and neutron detection devices for inverse kinematics, especially in combination with RI beams, has opened new horizons in nuclear physics. Research is progressing with detection devices optimized to explore these new and challenging area of physics. While some of the new detection developments aim for high energy and angular resolution, others focus on the increasing detection efficiency or enhancing large angular acceptance. Notable new detection systems such as STRASSE, TOGAXI, HYPATIA, NEBULA(plus), and TRIP have been developed at the RIBF to meet these demands. As high-intensity RI beams become available worldwide, we reflect on past detectors and provide a review of the future development of the detection devices.
@@ 1 Introduction
The Radioactive Isotope Beam Factory (RIBF) at the RIKEN Nishina Center utilizes a variety of advanced detection systems, including DALI2 takeuchi2014dali2 and NEBULA Kondo_NIM_2020 , which play a crucial role in exploring the structure and behavior of atomic nuclei. Additionally, external devices such as MINOS Obertelli_EPJA_2014 are often employed in collaborative experiments to enhance the facility’s experimental capabilities.
MINOS utilizes a cylindrical liquid hydrogen (LH2) target, with a maximum length of 150 mm along the beam axis, designed to enhance luminosity. This setup is particularly useful for studying exotic species produced at low count rates. Surrounding the target is a time projection chamber (TPC), which plays a key role in reconstructing the reaction vertex and identifying the reaction process. DALI2, an array of NaI(Tl) detectors, is dedicated to gamma-ray spectroscopy, offering high efficiency in detecting photons emitted from excited nuclear states. This capability is essential for nuclear structure studies. NEBULA, a neutron detection system, is optimized for time-of-flight measurements and can simultaneously detect multiple neutrons, enabling detailed studies of neutron-rich nuclei.
These detection systems has led to significant advancements in nuclear physics. Notable achievements include the discovery of new magic numbers Steppenbeck_Nature_2013 ; Taniuchi_Nature_2019 , the observation of a four-neutron resonance Duer_Nature_2022 , and the first spectroscopy of isotopes near or beyond the drip line Crawford_PhysRevLett_122_2019 ; Kondo_Nature_2023 . These discoveries highlight the importance of continuous development in detection systems, improving sensitivity and addressing the challenges posed by higher beam intensities expected in future experiments.
At RIBF, stable primary beams are accelerated up to 345 MeV/u using a set of four cyclotrons. The last two accelerators and the following beam-lines are shown Fig. 1. These beams are used to generate radioactive isotopes by fragmentation or fission at the primary target located at the F0 focal plane. These processes are highly effective for producing a wide range of neutron-rich and proton-rich species, as they involve high-energy collisions that break nuclei apart or induce asymmetric splitting. The BigRIPS separator Kubo_PTEP_2012 selects and identifies isotopes of interest on an event-by-event basis. These isotopes are then directed to one of three magnetic spectrometers: ZeroDegree Kubo_PTEP_2012 , SAMURAI Kobayashi_NIM_317_2013 , or SHARAQ Uesaka_NIM_266_2008 .
MINOS has been used at the F8 focal plane in combination with ZeroDegree and at the F13 focal plane coupled with SAMURAI. At F8, the focus is on high-efficiency gamma-ray detection using the DALI2 and DALI2+ Murray_RAPR_2018 arrays surrounding MINOS. This setup allows for the determination of bound nuclear excitation energies and decay patterns, providing valuable insights into nuclear structure.
Figure 1: RIBF beamline, equipments and the location of detectors.
The SAMURAI spectrometer is primarily designed for invariant mass spectroscopy, a technique central to the study of exotic nuclei and their decay modes. While missing mass spectroscopy in inverse kinematics was not a primary focus at SAMURAI in the past, the spectrometer is now capable of combining both invariant mass and missing mass spectroscopy through the integration of new detection systems. This combination allows for a more comprehensive understanding of complex reaction processes and facilitates a detailed analysis of nuclear structure. Additionally, gamma-ray spectroscopy has been conducted at F13. This manuscript reviews the current detection devices at RIBF for nuclear reactions and discusses upcoming upgrades and developments. Section 2 focuses on developments related to the detection of charged light particles and clusters, Section 3 summarizes the gamma-ray arrays used at RIBF and their future plans, while Section 4 addresses the neutron detection systems.
@@ 2 Missing-mass spectroscopy in inverse kinematics
The RIBF produces RI beams with energies suitable for conducting quasi-free knockout experiments. The advent of RIBF-stationed telescope arrays is essential for advancing missing-mass spectroscopy experiments. Although many physical interests have been explored, missing-mass spectroscopy using emitted particles from proton-induced knockout reactions could not realized with the MINOS device due to its limited angular resolution.
Missing masses $M_{miss}$ in inverse-kinematics proton-induced $(p,pX)$ knockout reactions are given by
$M_{miss}=\sqrt{\left\{E_{b}+m_{p}-(E_{p}+E_{X})\right\}^{2}-\left\{\vec{P_{b}}-(\vec{P_{p}}+\vec{P_{X}})\right\}^{2}}.$ (1)
@@
Figure 2: Relation between angles and energies of scattered particles X (left) and recoil protons (right) in inverse $p-X$ scattering.
Quasi-free elastic scattering reactions can be described using relatively simple models. The nucleons in an atomic nucleus can be assumed to follow the Fermi gas model, which treats the nucleus as a system with independent nucleons. The maximum wave number among the occupied states, the Fermi wave number, corresponds to the standard nuclear density of 0.17 fm-3. This value translates to $k=1.36~{}\textrm{fm}^{-1}$, or approximately 270 MeV/$c$ in the momentum space. When the momenta of the particles significantly exceed Fermi momentum due to reactions, the particles enter a quasi-free condition, no longer bound by the nuclear force. Achieving this condition requires incident beam energies to be sufficiently high. In inverse kinematics, the target proton gains momentum during the reaction, and that is nothing but the transferred momentum. Under the quasi-free approximation, the knocked-out particle is assumed to gain the same amount of momentum relative to the residual atomic nucleus, as illustrated in Fig. 3.
Figure 3: Left: Dependence of recoil proton momentum (wave-number) on laboratory angle. Right: Transfer momentum dependence of the differential cross section of $p-X$ scattering at 200 MeV of incident protons ermisch2005systematic ; hasell1986elastic ; moss1980proton . (right).
As the transferred momentum increases, the impact parameter between protons and intra-nuclear particles before the knockout becomes smaller. Consequently, the quasi-free scattering cross-sections tend to decrease. The right panel of Fig. 3 illustrates the relationship between cross-sections and momentum transfers. When designing experimental devices and setups, it is essential to balance the acceptance of distributions after quasi-free scatterings with the need to maximize cross-sections to achieve realistic yields.
2.1 The telescope array for $(p,2p)$ and $(p,3p)$ reactions
The recoil protons are emitted at wide angles in the laboratory frame for the $(p,2p)$ and $(p,3p)$ reactions. STRASSE (Silicon Tracker for RAdioactive nuclei Studies at SAMURAI Experiments) strasse_epja is a detection system under construction and optimized for $(p,2p)$ and $(p,3p)$ measurements at 200–250 MeV/nucleon at the RIBF facility. It consists of a compact silicon tracker placed in vacuum and a thick liquid hydrogen target with an effective diameter of 20 mm and a maximum length of 150 mm. STRASSE allows an optimum luminosity for beam intensities available at the RIBF. Thanks to the excellent reaction vertex resolution of 0.5 mm (FWHM), decent missing-mass spectra, which could not be achieved with the MINOS system, can be obtained with a resolution below 2 MeV. For $(p,2p)$ reactions at 250 MeV/nucleon, the one- and two-proton detection efficiencies of the silicon tracker are expected to be 86% and 49%, respectively. It is also designed to be used with large scintillator or germanium arrays.
@@ 2.1 The telescope array for $(p,2p)$ and $(p,3p)$ reactions
As shown in Fig. 4, the STRASSE silicon tracker is composed of two layers of double-sided silicon strip detectors (DSSDs) configured in a hexagonal shape. Each segment includes two DSSDs with strips parallel to the beam axis daisy chained. The inner segment has an active area of 30 mm $\times$ 244 mm with a thickness of 200 $\mu$m and a pitch size of 200 $\mu$m, while the outer segment has an active area of 62.6 mm $\times$ 242 mm with a thickness of 300 $\mu$m and a pitch size of 200 $\mu$m. The whole silicon tracker has a total of 17,358 electronic channels. 128-channel STS-XYTER sts-xyter ASIC chips with the ability to detect minimum ionizing particles are used to read out the signals from DSSDs via custom-made low-mass and low-capacitance microcables. Each STS-XYTER has a 14-bit time stamp and a 5-bit flash ADC. They are wire-bonded to the front-end PCBs (FEBs) placed in the radial direction inside the vacuum chamber to achieve the required ENC $\sim$ 10 keV. A cooling system is designed to dissipate the heat load of the electronics. Two star-shaped stainless steel frames support the detectors. The data from different FEBs are then sent to a concentration board called GBTxEMU via a mezzanine connection board. A PCI-express GERI board receives the aggregated data via optical fibers and synchronizes different GBTxEMUs.
The geometry of the LH2 target in STRASSE is designed to minimize the energy loss and angular straggling of the protons produced from quasi-free scattering. The target system consists of three parts: 1) the target cell made of 175 $\mu$m Mylar foil with a 31-mm cylinder and a 20-mm entrance window, 2) the cryostat to liquefy the hydrogen at $\sim$20 K, and 3) the control command system and the gas circuit to ensure the safe operation of the target. The hydrogen circulation in the STRASSE target is driven by the thermal siphon as the former PRESPEC prespec and MINOS Obertelli_EPJA_2014 targets. As shown in Fig. 4, two pumping stations located upstream and downstream of the STRASSE reaction chamber were used to achieve the required vacuum of the LH2 target. STRASSE will be a standalone system with windows to isolate from the beamline pipes.
STRASSE will be mainly used together with the CATANA (Cesium iodide Array for $\gamma$-ray Transitions in Atomic Nuclei at high isospin Asymmetry) array catana . It consists of CsI(Na) scintillators aligned at a radius of $\sim 20~{}\textrm{cm}$ to measure the energies of $\gamma$ rays and recoil protons with the SAMURAI setup. After the upgrade, CATANA now has 160 CsI(Na) crystals arranged into 8 rings (L1-L8) covering the polar angle from 17 to 77 degrees. For crystals at L2 to L8, Hamamatsu R580 PMTs with modified circuits read out the signals from the anode for $\gamma$ and the second last dynode for protons. Hamamatsu R11265 PMTs are used for crystals at L1. Note that the 40 Crystals at L1 and L2 are used only to detect protons. While for $\gamma$-ray detection, the energy resolution was measured to be 9% (FWHM) at 662 keV, for protons, the energy resolution was measured to be 1.7% (FWHM) using proton-proton elastic scattering at 230 MeV.
2.2 Elucidating cluster formation with $(p,pX)$ telescope
The particles released in knockout reactions are not necessarily to be protons and neutrons. By measuring cluster knockout reactions, it is possible to investigate the existence of clusters on the surface and inside of atomic nuclei. Alpha knockout reactions from stable tin isotopes, which is a stable nucleus, demonstrated the cluster formation in heavy nuclei tanaka2021formation . Analyzing the emitted particles after the $(p,p\alpha)$ reaction using a magnetic spectrometer, the cluster separation energy, a characteristic physical quantity of cluster knockout reaction, was deduced. It suggested that the isotopic dependence of cluster knockout reaction cross sections corresponds to the trend of alpha cluster formation predicted by theory. To extend the study towards the nuclei apart from the stability line, several inverse-kinematic alpha-knockout reactions in unstable nuclei experiments have been performed using solid and liquid hydrogen targets. In the following paragraphs, recent achievements of the particle spectroscopies performed in inverse kinematics are explained.
Figure 5: Experimental setup for 10Be$(p,p\alpha)$ reaction experiment. The figure is taken from li2023validation .
Figure 5 shows the experimental setup for measuring the alpha knockout reaction from a neutron-rich 10Be isotope. In this experiment, a 2-mm thick solid hydrogen target matsuda2011large was used. The surrounding detectors were placed far away so that the opening angle of the target thickness seen from the detector becomes small resulting in the alpha separation energy resolution of 1.06 MeV in sigma. As the emitted particles, protons, and alpha particles differ significantly in emission angle and energy, different detectors were placed at the respective emission angles. The recoil proton spectrometer (RPS-L(R)) chebotaryov2018proton ; matsuda2013elastic consisting of a drift chamber and a NaI (Tl) scintillator was prepared for the detection of the protons at 60 degrees from the beam axis. A silicon detector and CsI (Tl) telescope (Tele-L(R)) verde2013farcos was prepared for the alpha particles in forward angles of 4-12 degrees. Furthermore, the 6He residues were analyzed by the SAMURAI spectrometer installed forward. As a result, the triple differential cross section concerning the energy of protons suggested an alpha cluster structure mediating the two alphas + two neutrons picture of 10Be.
Figure 6: Silicon trackers for 8He$(p,p\alpha)$ reaction experiment. The figure is taken from Duer_Nature_2022 .
The alpha knockout reaction can also be used to reach uninvestigated nuclear systems. Here, the detection apparatus employed in the experiment for tetra-neutron search via 8He$(p,p\alpha)$ is introduced. As described in Fig. 6, silicon-strip detectors were placed downstream of the 50-mm-thick MINOS liquid hydrogen target. The detector arrangement was specialized for such events that alpha clusters in 8He nuclei collide with protons with small impact factors. The forward ejected protons and trailing alpha particles were tracked with three layers of strip silicon detectors, and their momenta were analyzed with the SAMURAI spectrometer.
In both cases described above, missing mass energy resolutions have been achieved as low as 1 MeV, however, the limitations in the solid angles are necessary to be improved. It will be crucial for the next-generation devices to inherit the successful concept of MINOS, thick target, and large solid angle, compensating for the weak RI beam intensity and achieving high detection efficiency. In addition, it would be ideal to measure multiple types of cluster knockout reactions over wide solid angles with one configuration. Incorporating a total energy calorimeter for the emitted particles can ensure the missing mass spectroscopy. Should the quasi-free scattering condition hold in inverse kinematics, these particles will be emitted at high energies towards relatively forward angles, while the recoil protons will be ejected to larger angles. The residues move forward along the beam axis. The detector array introduced in the following paragraph is designed to measure these charged particles at once, providing almost 1 MeV (sigma) missing mass energy resolution.
Figure 7: Graphical design of TOGAXSI array.

### v3

@@ Abstract
The advent of high-intensity radioactive ion beams has opened new avenues for nuclear structure research. By studying exotic ions, phenomena such as shell evolution, halos, and the limits of stability have been studied. In particular, Quasi-Free Scattering (QFS) experiments on hydrogen targets have proven to be a valuable tool to investigate the structure of exotic ions. Recently, a series of QFS experiments were performed at the Radioactive Isotope Beam Factory (RIBF) of the RIKEN Nishina Center employing the MINOS liquid hydrogen system. A fundamental part of the success of these experiments was the use of dedicated devices to measure the $\gamma$-rays, charged particles, and neutrons emitted in the reactions. The experience gained during the past campaigns, as well as the upcoming upgrade of the RIBF facility call for improvements on the existing devices, as well as for the development of new detection systems. Here we review the main detection devices used at the RIBF for QFS experiments, and give an overview of the ongoing and upcoming developments.
@@ 1 Introduction
Recent advances in the production of Radioactive Ion Beams (RIBs) have transformed modern nuclear physics by allowing us to study the structure of exotic nuclei. This has provided us essential tests for models of the nuclear force and of nuclear reactions under extreme conditions Nowacki_PPNP_120_2021 . Furthermore, precision measurements of nuclear masses, half-lives, and reaction cross sections have improved astrophysical models of supernovae and neutron star mergers Aumann_PPNP_112_2020 . At the Radioactive Isotope Beam Factory (RIBF) of the RIKEN Nishina Center Yano_NIMB2007 , the unparalleled high primary beam intensities Accelerator2021_RAPR ; Accelerator2022_RAPR have been exploited to study phenomena such as the limits of nuclear existence Kondo_Nature_2023 ; Ahn_PhysRevLett.129.212502_2022 , shell evolution Steppenbeck_Nature_2013 ; Taniuchi_Nature_2019 , and multi-nucleon resonances Kisamori_PhysRevLett_116_2016 ; Duer_Nature_2022 .
Figure 1 shows the layout of the RIBF. Stable beams are accelerated up to 345 MeV/u using four cyclotrons. The last two, the Intermediate-stage Ring Cyclotron (IRC), and the Superconducting Ring Cyclotron (SRC), can be seen in Fig. 1. After exiting the SRC, beams are delivered to the F0 focal plane of the BigRIPS fragment separator Kubo_PTEP_2012 where RIBs are produced via in-flight fragmentation or fission on a thick target. Reaction products are then separated and identified event-by-event based on their mass-to-charge ratio, $A/Q$, and atomic number, $Z$ Fukuda_NIMA2013 .
Figure 1: Schematic of the RIBF facility. After the SRC, stable beam are delivered to the BigRIPS fragment separator where RIBs are produced, separated and identified. The different spectrometers available at the RIBF are shown, and different devices used for quasi-free scattering experiments are listed in red, at the focal plane where they are placed.
The resulting secondary beams, with energies from a few tens to several hundreds MeV/u, are delivered to one of the spectrometers of the RIBF: ZeroDegree Kubo_PTEP_2012 , SAMURAI Kobayashi_NIM_317_2013 , SHARAQ Michimasa_NIMA_317_2013 , OEDO Michimasa_PTEP_2019 , and the Rare-RI Ring YAMAGUCHI2013 , where different types of measurement can be performed. In particular, Quasi-Free Scattering (QFS) experiments have been mainly performed at the ZeroDegree and SAMURAI spectrometers. In these measurements RIBs with energies around 200 MeV/u impinge on a secondary target, leading to the knockout of one or few nucleons and approximating a free nucleon-nucleon collision. With this technique bound and unbound excited states, single-particle properties, and correlations within nuclei can be probed.
In recent years, a series of experiments involving QFS were performed at the RIBF employing the MINOS liquid hydrogen target Obertelli_EPJA_2014 ; santamaria2018tracking . Thanks to the time projection chamber (TPC) surrounding the target, it was possible to deduce the interaction vertex with a resolution of about 5 mm (FWHM) which, in turn, allowed for the use of a thicker target with higher luminosity. Different experiments with target lengths varying between 10 and 15 cm were performed at F8 and F13, and different detection systems were employed to measure the reaction residues.
At F8, the main focus was in-beam $\gamma$-ray spectroscopy following knockout reactions. In this case the DALI2+ array takeuchi2014dali2 ; Murray_RAPR_2018 was employed. This array, consisting of 226 NaI(Tl) detectors, provides high $\gamma$-ray detection efficiency which enables the study of bound excited states, as well as exclusive cross sections. At F13, MINOS was followed by the SAMURAI spectrometer to perform invariant mass measurements. In this case, in addition to the standard detectors of SAMURAI, which measure the $A/Q$ and $Z$ of the outgoing fragments, the neutron detectors NeuLAND Boretzky_NIM_2021 and NEBULA Kondo_NIM_2020 were employed. Furthermore, the DALI2+ array was also used in some experiments at F13 to measure excited states.
The combination of MINOS with these different detection systems was successfully used to measure phenomena such as magicity around 78Ni Taniuchi_PTEP_2025 , shell evolution around the Ca isotopes Lui_PTEP_2025 and the $N=40$ Island of Inversion (IoI) Cortes_PTEP_2025 , the limits of the $N=20$ IoI Kahlbow_PTEP_2025 , and the deformation of isotopes between Zn and Zr Chen_PTEP_2025 . These measurements also highlighted possible improvements to the detections systems and provided new ideas to further optimize the measurements. The continuous developments on the instrumentation is, therefore, of fundamental importance for studies involving QFS. This is particular the case at the RIBF, where an upgrade plan of the facility has been outlined for the coming decade ribf_upgrade .
In this article we review the main detection systems used for QFS experiments at the RIBF and give an overview of the on-going and future developments. Section 2 focuses on charged particle detectors for missing mass measurements, in particular STRASSE and TOGAXI. Section 3 gives an overview on the developments for in-beam $\gamma$-ray spectroscopy such as the HYPATIA array. Section 4 focuses on neutron detection arrays, in particular HIME and NEBULA+. In Sec. 5 a summary and outlook are presented.
@@ 2 Missing-mass spectroscopy in inverse kinematics
The nucleons in the atomic nucleus can be considered to be a system of independent particles within the Fermi gas model. The Fermi wave number, which defines the momentum of the highest occupied state in nuclear matter, is related to the typical nuclear saturation density of $\rho_{0}\approx 0.17~\mathrm{fm}^{-3}$. This value translates to $k_{F}=[3\pi^{2}/(2\rho_{0})]^{1/3}=1.36~\textrm{fm}^{-1}$, or approximately 270 MeV/$c$ in the momentum space. When the momenta of the particles significantly exceed Fermi momentum, the particles enter a quasi-free condition, no longer bound by the nuclear force. Therefore, proton-induced knockout reactions in inverse kinematics, expressed as $(p,pX)$, at beam energies above 200 MeV/u can be considered QFS reactions, where the proton scatters off a quasi-free cluster or nucleon inside the target nucleus.
((a))
((b))
Figure 2: (a) Dependence of recoil proton momentum on the scattering angle for different knockout reactions at 200 MeV/u. (b) Dependence of the QFS differential cross section on the momentum transfer ermisch2005systematic ; hasell1986elastic ; moss1980proton .
In $(p,pX)$ reactions the proton, initially at rest in the laboratory frame, gains momentum due to the reaction with the incoming heavy ion. Under the QFS approximation, it is assumed that the knocked-out particle acquires the same momentum transfer relative to the residual nucleus, as if the reaction was a direct, two-body collision, largely unaffected by the rest of the nucleus. Figure  2a) illustrates the dependence of the momentum transferred to the recoil proton as a function of the scattering angle. As the transferred momentum increases, the impact parameter between protons and intra-nuclear particles before the knockout becomes smaller. Consequently, the quasi-free scattering cross-sections tend to decrease, as shown in Fig. 2b).
Missing-mass spectroscopy following QFS reactions in inverse kinematics has become a key technique for investigating exotic nuclear structures, particularly in weakly bound and unbound nuclei, clustered systems, or exotic decays where direct detection is impractical. In this technique, undetected particles are identified by applying conservation of energy and momentum. Instead of detecting all final-state particles directly, the momenta and energies of the incoming beam and some of the outgoing particles are measured. The mass of the unobserved particle(s) is then deduced as the “missing” component needed to balance the reaction kinematics.
In particular for $(p,pX)$ reactions, the missing mass, $M_{\text{miss}}$, is deduced as
$M_{\text{miss}}=\sqrt{\left\{E_{b}+m_{p}-(E_{p}+E_{X})\right\}^{2}-\left\{\vec{P_{b}}-(\vec{P_{p}}+\vec{P_{X}})\right\}^{2}},$ (1)
where $E_{i}$, and $\vec{P_{i}}$ represent the energy and the three-momentum vector of the beam ($i=b$), the recoil proton ($i=p$) and the knocked-out particle ($i=X$), respectively, and $m_{p}$ is the proton mass.
((a))
((b))
Figure 3: Relation between emission angle and energy of (a) the scattered particles X, and (b) the recoil protons in $(p,pX)$ reactions at 200 MeV/u. The reaction Q-values are assumed to be zero for these calculations.
Figure 3 shows the relation between the energy, and emission angle, $\theta$, of a) the knocked-out particles (such as protons, deuterons, 3He, tritons, and $\alpha$ particles), and b) the recoil protons, for $(p,pX)$ reactions at incident energy of 200 MeV/u. For simplicity, the internal momentum of the knocked-out particle in the initial state is not taken into account. It can be seen that the mass of the knocked-out particles significantly influences $\theta_{X}$, particularly the maximum angles in which detection is possible: While knocked out protons can be detected at a larger angular range, heavier ejectiles are only emitted at forward angles, smaller than 30. In contrast, as can be seen on Fig. 3b), the recoil protons are scattered up to 90, regardless of the knocked-out particle.
When designing instrumentation for QFS reactions, it is essential to consider the angular distributions and energies of the scattered particles, as well as the cross sections required to achieve realistic yields. Detectors with different characteristics are therefore necessary to address different physics questions that can be studied with QFS reactions. The next subsections describe two major ongoing developments: STRASSE, optimized for nucleon knockout, and TOGAXSI, designed to study cluster knockout reactions. Each system is introduced with its specific motivation and improvement over existing approaches.
2.1 STRASSE: A telescope array for $(p,2p)$ and $(p,3p)$ reactions
STRASSE (Silicon Tracker for RAdioactive nuclei Studies at SAMURAI Experiments) strasse_epja is a detection system optimized for $(p,2p)$ and $(p,3p)$ reactions at 200–250 MeV/nucleon. Similarly to MINOS, a thick LH2 target with a length up to 150 mm is used to induce the QFS reactions. The recoil protons are detected, not by a TPC, but by a highly segmented Si tracker array placed in vacuum to minimize the angular straggling of protons. With this setup, proton detection efficiencies of 86% and 49% for one and two protons, respectively, can be achieved, comparable to those obtained with MINOS. Thanks to the high segmentation of the Si tracker, a vertex resolution of 0.5 mm (FWHM) and an angular resolution of $\sim 12$ mrad (FWHM), can be obtained.
STRASSE will be coupled with the CATANA (Cesium iodide Array for $\gamma$-ray Transitions in Atomic Nuclei at high isospin Asymmetry) array catana , composed of 140 CsI(Na) scintillators arranged into 7 rings (L2-L8), and covering polar angles from 17 to $7$. CATANA crystals can stop protons up to 250 MeV. In particular, a dual-gain readout has been developed for CATANA to enable simultaneous $\gamma$-ray and proton measurement. Note that the 20 crystals in L2 will be used to detect only protons. With the STRASSE + CATANA setup, missing-mass spectra, which could not be achieved with the MINOS system, are expected to be available with a resolution better than 4 MeV (FWHM). This offers an excellent opportunity to study highly unbound states (such as 4n or 6n states) or isomeric bound states (such as the $0_{2}^{+}$ states of doubly magic nuclei) at the RIBF.
@@ 2.1 STRASSE: A telescope array for $(p,2p)$ and $(p,3p)$ reactions
2.2 TOGAXI: Elucidating cluster formation with $(p,pX)$ telescope
Clustering is a fundamental feature of nuclear structure, where nucleons tend to group into substructures such as $\alpha$ particles within a nucleus. By measuring cluster knockout reactions, it is possible to investigate the existence of clusters on the surface and inside of atomic nuclei. A notable example of $\alpha$-cluster formation in heavy ions was recently reported for the stable Sn isotopes Tanaka_Science_2021 . Analyzing the emitted particles after the $(p,p\alpha)$ reaction using a magnetic spectrometer, the cluster separation energy, a characteristic physical quantity of cluster knockout reaction, was deduced. Based on its agreement with theoretical models, a close connection between the $\alpha$-cluster formation and the neutron skin was suggested.
Similar types of studies have also been performed for exotic isotopes by using liquid or solid hydrogen targets and inverse kinematics. For example, in a recent measurement at the SAMURAI spectrometer, the measured triple differential cross section as a function of the proton energy revealed the existence of $\alpha$-clusters in 10Be, in agreement with a $\alpha-2$$n$ configuration matsuda2011large . Figure 5 shows the experimental setup used in that study. A 2-mm-thick solid hydrogen target (SHT) matsuda2011large was employed to induce $(p,p\alpha)$ reactions. The 6He residual nuclei were analyzed using the forward-located SAMURAI spectrometer. Because the emitted protons and alpha particles differ significantly in both emission angle and energy, separate detectors were placed to match each particle’s emission characteristics. Protons emitted at 60 degrees relative to the beam axis were measured with a recoil proton spectrometer (RPS-L(R)), composed of a drift chamber and a NaI (Tl) scintillator chebotaryov2018proton ; matsuda2013elastic , and $\alpha$ particles emitted between 4 and 12 degrees, were detected using a telescope consisting of a silicon detector and a CsI (Tl) scintillator (Tele-L(R)) verde2013farcos . By placing the detectors at a relatively large distance from the target, which minimizes the relative opening angle, an $\alpha$ separation energy resolution of 2.5 MeV (FHWM) was achieved. In addition, the time of flight of the emitted protons was measured.
[…]
In the cases described above, missing mass energy resolutions as low as 2.4 MeV (FWHM) has been achieved; however, the limited solid angle coverage remains a point for improvement. For next-generation detectors, it will be essential to inherit the successful concepts demonstrated by MINOS – namely, the use of a thick target and a detection system with large solid-angle coverage – to compensate for the low intensity of RI beams and to achieve high detection efficiency. Ideally, the system should be capable of measuring multiple types of cluster knockout reactions over a wide solid angle within a single configuration. Incorporating a total energy calorimeter for the emitted particles can further ensure accurate missing mass spectroscopy. The detector array introduced in the following paragraph is designed to simultaneously detect these charged particles, enabling a missing mass energy resolution of approximately 3 MeV (FWHM).
Figure 7 shows the TOGAXSI (TOtal energy measurement by GAgg and verteX measurement by SIlicon strips) telescope TANAKA20234 , being developed within the “ONOKORO” project uesaka2021comprehensive for dedicated measurements of cluster knockout reaction at the RIBF. Similarly to MINOS, it uses a thick LH2 target which provides high luminosity. For ONOKORO, the same type of liquid hydrogen target used for the STRASSE has been adopted. The target is surrounded by a Si tracker for vertex reconstruction and a GAGG:Ce calorimeter for energy measurement, which ensures accurate missing mass spectroscopy.
Figure 7: Graphical design of TOGAXSI array. Incoming beams interact in the LH2 target to produce cluster-knockout reactions. A Si tracker and a GAGG:Ce calorimeter are employed to measure the scattered clusters and the recoil protons and performe missing mass measurements.
[…]
As can be seen in Fig. 3a), under QFS conditions in inverse kinematics, the emitted clusters have high energies and are directed towards relatively forward angles, below 30, whereas recoil protons are emitted at larger angles. Consequently, the detector is designed to detect clusters at polar angles of $8^{\circ}$$30^{\circ}$. For the detection of recoil protons, detection angles between $35^{\circ}$$70^{\circ}$ are selected. As can be seen in Fig. 2, for these angles a momentum transfer between 2233 fm-1 is expected. This in turn leads to cross sections from $10^{-1}$ to $10^{-2}$ mb/Sr, which allows to keep a realistic yield for the experiments. In these angles the expected energy of the protons varies from 8080250250 MeV, and the cluster energy is between 7575180180 MeV/u.
The construction of the full array is underway. TOGAXSI is expected to accelerate comprehensive cluster research with missing mass spectroscopy based on the detection of charged particles. Table 1 summarizes the properties of MINOS, STRASSE and TOGAXI.
Table 1: Comparison of inverse-kinematics proton-induced reaction detector systems at RIBF.
Parameter MINOS Obertelli_EPJA_2014 STRASSE strasse_epja TOGAXSI TANAKA20234
Vertex resolution (FWHM) $\sim 5$ $m$ $\sim 0.5$ $m$ $\sim 0.5$ $m$
Angular resolution (FWHM) $\sim 38$ mrad $\sim 12$ mrad $\sim 10$ mrad
Missing-mass resolution (FWHM) $\leq 4$ MeV $\sim 3$ MeV
Angular coverage (cluster) $8^{\circ}$$30^{\circ}$
Angular coverage (recoil $p$) $14^{\circ}$$70^{\circ}$ $14^{\circ}$$70^{\circ}$ $35^{\circ}$$70^{\circ}$